The perimeter of a triangle is 47km. the first side is 5km shorter than twice the second. The third side is 2km longer than the first find the length of all sides
Answer by rfer(16322)

1st side=2x-5=17
2nd side=x=11
3rd side=2x-3=19
(2x-5)+x+(2x-3)=47
5x-8=47
5x=55
x=11